# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v2

@@ 4.1 Hierarchical Data Representation
For hyper-relational knowledge graphs, we regard key-value pairs as complementary information for facts. Thus, we translate H-Facts $\mathcal{F}_{HKG}$ =$\text{=}\,\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{HiDR}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{HKG}\}$ following the definition, where $\mathcal{F}_{c}\text{=}\{(f,has\,relation,e_{r}),(f,has$ $,\,head\,entity,h),(f,has\,tail\,entity,t),(f,k_{1},v_{1}),$$\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,((h,r,t),\{(k_{i}\text{:}\,$ $v_{i})\}_{i=1}^{m})\in\mathcal{F}_{HKG})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
For nested factual knowledge graphs, HiDR can naturally represent hierarchical facts, so we translate the atomic facts $\mathcal{F}_{NKG}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts ^$\hat{\mathcal{F}}_{NKG}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},$ $r_{2},t_{2}))|(h_{i},r_{i},t_{i})\in\mathcal{F}_{NKG}\}$ into the form of HiDR that $\mathcal{G}^{HiDR}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{NKG}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})\,|\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{NKG}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},has$ $tail\,entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},$ $t_{i})\in\mathcal{F}_{NKG}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{NKG}\}$.
For temporal knowledge graphs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes. Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{TKG}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{TKG}^{HKG}\text{=}\,\{(h,r,t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$. Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{TKG}^{HiDR}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}_{TKG}^{HiDR}\}$ following the definition,where $\mathcal{F}_{a}\,\text{=}\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,$ $has\,head\,entity,h),\,(f,has\,tail\,entity,t),\,(f,$ $begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\,\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
@@ 4.2 Hierarchical Structure Learning
In this section, we illustrate how various KGs in the form of HiDR can be encoded by the Hierarchical Structure Learning module, abbreviated as HiSL shown in Fig. 3.
Figure 3: HiSL module for intra-fact and inter-fact MP.
@@
We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{HiDR}|\times d}$ for atomic nodes and all relation edges. Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$, define as:
$\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}},$ (1)
where $\mathbf{E_{a}}\subseteq\mathbf{E},\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$ denote the atomic relation embeddings and a projection matrix. Then we initialize the fact node embeddings $\mathbf{H}_{f}$ to explicitly capture key information within facts by utilizing the embedding of (main) triple:
@@
where $(h,r,t)\in\mathcal{F}_{a}$, the operation [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the vector concatenation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding and $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP.
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we utilize the Time2Vec Kazemi et al. (2019) to encode the timestamp $\tau$ into an embedding:
@@
In this stage, massage passing is conducted for fact nodes. Given a fact node $f_{k}\in\mathcal{V}_{f}$, we construct its constituent elements, i.e., one-hop neighbors, as the node set $\mathcal{V}_{k}\,\text{=}\,\{v\in\mathcal{N}_{f_{k}}\,|\,v\in\mathcal{V}_{a}\cup\mathcal{V}_{r}\}$, where $\mathcal{N}_{f_{k}}$ is the set of one-hop neighbors of fact node $f_{k}$. Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{HiDR}$. For this subgraph, we employ the graph attention network (GAT) Brody et al. (2021) to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$. The formula for calculating $\alpha_{i,j}$ in the $l$-th layer is as follows:
@@
$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$
@@
At this stage, message passing is conducted on the whole graph $\mathcal{G}^{HiDR}$. Similar to previous work Vashishth et al. (2019), we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges. Specifically, we employ the circular-correlation operator inspired from HolE Nickel et al. (2016), defined as:
$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right)$
where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, the $\odot$ is the element-wise (Hadamard) product. Furthermore, in order to fully capture the heterogeneity of the graph, we classify edges along two dimensions: $\lambda(r)\in\left\{forward,reverse\right\}$ and $\tau(r)\in\{connected\,\,relation,atomic\,\,relation,nested\,$ $relation\}$ and adopt two relation-type specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for more fine-grained aggregation:
$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l}$
$\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l}$
where $\mathbf{W}^{l}_{self},\mathbf{W}^{l}_{rel}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid activation function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$. We utilize $\phi\left(\cdot\right)$ to combine the information from both edge $r$ and node $j$, and then passes it to node $i$. Subsequently, node $i$ aggregates the information according to the types of edge $r$ separately to update its embedding, while edge $r$ is also projected into the same embedding space as the updated nodes.
@@ 4.3 Link Prediction Decoder
Since the query varies across different settings, we use the transformer Vaswani et al. (2017) as the decoder with the mask pattern. Specifically, we convert the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input to the transformer. Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss function:
$P=\,\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right),$
$\mathcal{L}=\sum_{t=0}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}{y_{t}\log P_{t}}$
@@
Table 1: Results of link prediction on HKG datasets. All baselines’ results are taken from Chung et al. (2023). The best results are written bold, while the second are underlined. # Tr. Params denotes the number of learnable parameters during training.
WikiPeople WD50K
Model # Tr. Params subject/object all entities # Tr. Params subject/object all entities
MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10 MRR Hits@1 Hits@10
NaLP - 0.356 0.271 0.499 0.360 0.275 0.503 - 0.230 0.170 0.347 0.251 0.187 0.375
tNaLP - 0.358 0.288 0.486 0.361 0.290 0.490 - 0.221 0.163 0.331 0.243 0.182 0.360
RAM 27.34M 0.459 0.384 0.584 0.461 0.386 0.585 - 0.276 0.210 0.399 0.296 0.232 0.416
HINGE - 0.393 0.309 0.547 0.395 0.311 0.549 - 0.264 0.187 0.410 0.277 0.200 0.424
NeuInfer - 0.357 0.247 0.533 0.357 0.248 0.532 - 0.220 0.154 0.347 0.225 0.158 0.355
StarE 7.84M 0.458 0.364 0.611 - - - 10.39M 0.309 0.234 0.452 - - -
HyTransformer - 0.460 0.382 0.594 - - - - 0.304 0.231 0.443 - - -
GRAN 15.26M 0.462 0.366 0.610 0.465 0.371 0.613 18.51M 0.330 0.255 0.472 0.361 0.286 0.501
HyNT 23.02M 0.482 0.415 0.602 0.481 0.414 0.603 29.61M 0.333 0.259 0.474 0.360 0.287 0.500
UniHR 8.02M 0.491 0.417 0.618 0.493 0.420 0.621 10.55M 0.348 0.278 0.482 0.382 0.313 0.513
Table 2: Results of base link prediction (left) and triple prediction (right). All baselines’ results are taken from Xiong et al. (2024). For BiVE Chung and Whang (2023) and NestE Xiong et al. (2024), we pick their best variants.
Model FBHE/FBH DBHE FBH FBHE DBHE
MRR Hits@10 MRR Hits@10 #Tr. Params MR MRR Hits@10 #Tr. Params MR MRR Hits@10 #Tr. Params MR MRR Hits@10
Base link prediction Triple prediction
QuatE 0.354 0.581 0.264 0.440 - 145603.8 0.103 0.114 - 94684.4 0.101 0.209 - 26485.0 0.157 0.179
BiQUE 0.356 0.583 0.274 0.446 - 81687.5 0.104 0.115 - 61015.2 0.135 0.205 - 19079.4 0.163 0.185
Neural-LP 0.315 0.486 0.233 0.357 - 115016.6 0.070 0.073 - 90000.4 0.238 0.274 - 21130.5 0.170 0.209
DRUM 0.317 0.490 0.237 0.359 - 115016.6 0.069 0.073 - 90000.3 0.261 0.274 - 21130.5 0.166 0.209
AnyBURL 0.310 0.526 0.220 0.364 - 108079.6 0.096 0.108 - 83136.8 0.191 0.252 - 20530.8 0.177 0.214
BiVE 0.370 0.607 0.274 0.422 12.66M 6.20 0.855 0.941 12.67M 8.35 0.711 0.866 10.74M 3.63 0.687 0.958
NestE 0.371 0.608 0.289 0.443 12.11M 3.34 0.922 0.982 12.46M 3.05 0.851 0.962 10.22M 2.07 0.862 0.984
UniHR 0.401 0.619 0.300 0.455 4.12M 2.46 0.946 0.993 4.12M 5.20 0.793 0.890 3.67M 1.90 0.862 0.987

### v3

@@ 4.1 Hierarchical Data Representation
For hyper-relational knowledge graphs, we regard key-value pairs as complementary information for facts. Thus, we translate the H-Facts $\mathcal{F}_{HKG}\,\text{=}$
$\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{HiDR}_{HKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{HKG}\}$ following the definition, where $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,has$ $\,head\,entity,h),(f,has$
$tail\,entity,t),(f,k_{1},v_{1}),\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})\in\mathcal{F}_{HKG})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
For nested factual knowledge graphs, HiDR can naturally represent hierarchical facts, so we translate the atomic facts $\mathcal{F}_{NKG}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts ^$\hat{\mathcal{F}}_{NKG}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},$ $r_{2},t_{2}))|(h_{i},r_{i},t_{i})\in\mathcal{F}_{NKG}\}$ into the form of HiDR that $\mathcal{G}^{HiDR}_{NKG}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{HiDR}_{NKG}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})\,|\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{NKG}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},has\,tail$
$entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},$ $t_{i})\in\mathcal{F}_{NKG}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{NKG}\}$.
For temporal knowledge graphs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes. Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{TKG}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{TKG}^{HKG}\text{=}\,\{(h,r,t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$. Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{TKG}^{HiDR}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}_{TKG}^{HiDR}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}$ $\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$, $\mathcal{F}_{c}\,\text{=}\,\{(f,has\,relation,e_{r}),(f,$ $has\,head\,entity,h),\,(f,has\,tail\,entity,t),\,(f,$ $begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,$
$begin\text{:}\,\tau_{b},end\text{:}\,\tau_{e})\in\mathcal{F}_{TKG}^{HKG}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
@@
Figure 3: HiSL module for intra-fact and inter-fact message passing.
@@ 4.2 Hierarchical Structure Learning
It’s evident that HiDR form introduces many additional relation nodes and fact nodes. To avoid significantly increasing the model’s training parameters while fully capturing the hierarchy of HiDR form, we design a Hierarchical Structure Learning module, abbreviated as HiSL shown in Fig. 3.
@@
We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{HiDR}|\times d}$ for atomic nodes and all relation edges. Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$ with a projection matrix $\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$: $\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}}$, where $\mathbf{E_{a}}\subseteq\mathbf{E}$ is the atomic relation embeddings. Then we initialize the fact node embeddings $\mathbf{H}_{f}$ to explicitly capture key information within facts by utilizing the embedding of (main) triple:
@@
where $(h,r,t)\in\mathcal{F}_{a}$, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the concatenation operation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding and $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP. Therefore, the initialization of relation nodes and fact nodes is sufficiently parameter-efficient.
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec [18]:
@@
In this stage, massage passing is conducted for fact nodes. Given a fact node $f_{k}\in\mathcal{V}_{f}$, we construct its constituent elements, i.e., one-hop neighbors, as the node set $\mathcal{V}_{k}\,\text{=}\,\{v\in\mathcal{N}_{f_{k}}\,|\,v\in\mathcal{V}_{a}\cup\mathcal{V}_{r}\}$, where $\mathcal{N}_{f_{k}}$ is the set of one-hop neighbors of fact node $f_{k}$. Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{HiDR}$. For this subgraph, we employ the graph attention network [5] to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$. The formula for calculating $\alpha_{i,j}$ in the $l$-th layer is as follows:
@@
$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$ (4)
@@
At this stage, message passing is conducted on the whole graph $\mathcal{G}^{HiDR}$. Similar to previous work [31], we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges. We employ the circular-correlation operator, defined as:
$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right),$ (5)
where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, the $\odot$ is the element-wise (Hadamard) product. Furthermore, in order to fully capture the heterogeneity of the graph, we classify edges along two dimensions: $\lambda(r)\in\left\{forward,reverse\right\}$ and $\tau(r)\in\{connected\,\,relation,atomic\,\,$
$relation,nested\,relation\}$ and adopt two relation-type specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for more fine-grained aggregation:
$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l},$ (6)
$\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l},$ (7)
where $\mathbf{W}^{l}_{self},\mathbf{W}^{l}_{rel}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid activation function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$. We utilize $\phi\left(\cdot\right)$ to combine the information from edge $r$ and node $j$, and then passes it to node $i$ for update.
Through Intra-fact and Inter-fact two-stage message passing, nodes can fully capture both local semantic and global structural information. Moreover, its number of training parameters does not increase with the scale of the graph, thereby effectively adapting to the HiDR form.
@@ 4.3 Link Prediction Decoder
Since the query varies across different settings, we use the transformer [32] as the decoder with the mask pattern. Specifically, we convert the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input to the transformer. Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss:
$P=\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right),$ (8)
$\mathcal{L}=\sum_{t=0}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}{y_{t}\log P_{t}},$ (9)